OpenStudy (anonymous):

How do you factor trinomials? Example: n2+ 6n + 5

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to 5 AND add to 6

OpenStudy (anonymous):

@jim_thompson5910 What two numbers multiply to 5 and add to 6?

jimthompson5910 (jim_thompson5910):

any ideas of what they could be?

OpenStudy (anonymous):

OH , 1 & 5? Haha. @jim_thompson5910 Blonde moment.

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

so n^2+6n + 5 factors to (n+1)(n+5)

OpenStudy (anonymous):

So then it's just (n+1)(n+5) correct? @jim_thompson5910 Thanks so much!

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

that's all there is to it

OpenStudy (anonymous):

That's easier than I thought. I appreciate it!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

What about if its an equation like n2+7n+10? Nothing computes to those?

jimthompson5910 (jim_thompson5910):

list out the factors of 10

OpenStudy (anonymous):

1,2,5,10?

jimthompson5910 (jim_thompson5910):

so we have 1+10 = 11 .... so 1 and 10 aren't it 2+5 = 7 ... that's it, we found it, so we're done notice how 1*10 = 10 and 2*5 = 10

OpenStudy (anonymous):

Oh! Awesome thanks! :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np